# Teacher Encoder-Student Decoder Denoising Guided Segmentation Network for Anomaly Detection

### v3

@@ 4.4. Results
Table 4. Ablation study on the modules: The results of RCM, AFF, and PCAR individually and in combination, evaluated on the MVTec AD dataset, for image-level AUC (%), pixel-level AP (%), and instance-level IAP (%).
@@
4.5. Ablation Studies
In Tab. 4, we conducted ablation experiments to effectively evaluate the proposed three modules. Specifically, we adopted the network architecture from (Zhang et al., 2022) as the baseline and performed a detailed analysis and comparative experiments to assess the independent effects of each module as well as their combined impact on network performance. The experimental results demonstrate that, by comparing Experiment 2 with 5 and Experiment 3 with 6, integrating the proposed PCAR module with either the RCM or AFF module significantly enhances the network’s performance. Furthermore, the results of Experiment 7 further reveal that incorporating the PCAR, RCM, and AFF modules into the network architecture achieves optimal performance. This combination not only enables more accurate detection and localization but also substantially improves the overall network effectiveness, thereby validating the efficiency and superiority of our proposed approach.
5. Conclusion
In this paper, we propose a novel PCAR module and integrate it into our PFADSeg model. The PCAR module employs parallel convolution techniques to effectively capture multi-scale spatial information, enabling precise localization and segmentation of anomalous regions while suppressing irrelevant noise in detection images. The PFADSeg model further optimizes the feature extraction process between the student and teacher networks, using an improved strategy to enhance discriminative features in anomalous regions. Additionally, a segmentation network is integrated to adaptively fuse features from the student-teacher network, achieving more accurate anomaly segmentation. Experiments on the MVTec AD dataset demonstrate that our method outperforms state-of-the-art approaches. Specifically, in image-level anomaly detection, our method achieves an AUC of $98.9\%$, surpassing the current best by $0.4\%$. For pixel-level anomaly localization, it achieves an AP of $76.4\%$, exceeding the best by $0.6\%$, with AUC nearing the state-of-the-art. Notably, for instance-level anomaly detection, our method improves IAP@$90$ by $7.3\%$ and IAP by $4.1\%$. These results validate the effectiveness and superiority of our approach.

### v4

@@ 4.5 Ablation Studies
Table 4: Ablation study on the modules: The results of RCM, AFF, and PCAR individually and in combination, evaluated on the MVTec AD dataset, for image-level AUC (%), pixel-level AP (%), and instance-level IAP (%).
@@
5 Conclusion
In this paper, we propose a novel PCAR module and integrate it into our PFADSeg model. The PCAR module employs parallel convolution techniques to effectively capture multi-scale spatial information, enabling precise localization and segmentation of anomalous regions while suppressing irrelevant noise in detection images. The PFADSeg model further optimizes the feature extraction process between the student and teacher networks, using an improved strategy to enhance discriminative features in anomalous regions. Additionally, a segmentation network is integrated to adaptively fuse features from the student-teacher network, achieving more accurate anomaly segmentation. Experiments on the MVTec AD dataset demonstrate that our method outperforms state-of-the-art approaches. Specifically, in image-level anomaly detection, our method achieves an AUC of $98.9\%$, surpassing the current best by $0.3\%$. For pixel-level anomaly localization, it achieves an AP of $76.4\%$, exceeding the best by $0.5\%$, with AUC nearing the state-of-the-art. Notably, for instance-level anomaly detection, our method improves IAP@$90$ by $7.3\%$ and IAP by $4.1\%$. These results validate the effectiveness and superiority of our approach.
Acknowledgments
This preprint has not undergone peer review (when applicable) or any post-submission improvements or corrections. The Version of Record of this contribution is published in Neural Information Processing, and is available online at https://doi.org/10.1007/978-981-95-4109-6_17.